# Auto Arena of LLMs: Automating LLM Evaluations with Agent Peer-battles and Committee Discussions

### v3

@@ Abstract
As LLMs evolve on a daily basis, there is an urgent need for a trustworthy evaluation method that can provide robust evaluation results in a timely fashion. Currently, as static benchmarks are prone to contamination concerns, users tend to trust human voting platforms, such as Chatbot Arena. However, human annotations require extensive manual efforts. To provide an automatic, robust, and trustworthy evaluation framework, we innovatively propose the Auto-Arena of LLMs, which automates the entire evaluation process with LLM agents. Firstly, an examiner LLM devises queries. Then, a pair of candidate LLMs engage in a multi-round peer-battle around the query, during which the LLM’s true performance gaps become visible. Finally, a committee of LLM judges collectively discuss and determine the winner, which alleviates bias and promotes fairness. In our extensive experiment on the 17 newest LLMs, Auto-Arena shows the highest correlation with human preferences, providing a promising alternative to human evaluation platforms.111  We release the code at https://github.com/DAMO-NLP-SG/Auto-Arena-LLMs.
@@ 1 Introduction
Since ChatGPT and GPT-4 [34] gained popularity, Large Language Models (LLMs) have risen to the forefront of technological innovation, capturing broad industry and social interests [40]. This enthusiasm has spurred numerous organizations to release their own LLMs [38, 37]. However, the rapid pace at which these models are released and updated poses a significant challenge for users attempting to understand their capabilities and monitor their evolution. Consequently, there has been a pressing demand for comprehensively evaluating LLMs recently [10].
@@
One line of these evaluation researches conducts automatic evaluation with static datasets. Among these, static datasets with predefined metrics, such as GSM8k [13] and MMLU [22], are constructed with aspect-specific input-output pairs, such as human exam-type questions and their corresponding answers. Given the questions, the LLM-produced answers are compared to ground-truth answers using metrics such as accuracy. However, these approaches suffer significantly from contamination concerns [35], where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing the evaluation results. The closed-formed ground-truth answers also limit their utility in assessing models’ performances on general or open-ended questions. Model-based evaluation, such as MT-Bench [44] and AlpacaEval [19], provides an alternative for evaluating LLMs on open-ended questions. These methods typically ask two models to generate responses to the same open-ended question and then employ a strong judge model (e.g., GPT-4) to choose the better response. However, the static question sets still pose the issue of contamination. Additionally, the assumption of the existence of a strong judge model makes the evaluation framework less generalizable.
@@
Aside from automated evaluations, human assessment, although requiring significant manual efforts, remains the gold standard for most users. A noticeable example is Chatbot Arena [44], a crowdsourced voting platform that gathers anonymous votes on LLM performances and calculates ELO scores to rank these models. The resulting leaderboard [^2] https://leaderboard.lmsys.org/ is widely considered a trustworthy indicator of an LLM’s general capabilities. However, a reliable model evaluation on this platform must be supported by a large number of human votes, which requires considerable time and effort. Consequently, when newly developed models enter the scene, they often struggle to quickly amass a large number of votes. Moreover, this strong reliance on human votes limits its application in various scenarios. For example, the performance of non-English languages is difficult to estimate, as most queries on the platform are in English. The completely open participation may also result in uneven evaluation quality.
Figure 1: An illustration of Auto Arena of LLMs (Auto-Arena).
@@
To provide automatic, reliable, and human-like LLM evaluations, we propose Auto Arena of LLMs (Auto-Arena), a framework that automates the entire LLM evaluation process with LLM agents. As shown in Figure 1, the Auto-Arena framework consists of three stages. Firstly, an LLM examiner agent is tasked with generating questions, mimicking real-life users posting queries. Secondly, two candidate LLMs interact with each other and engage in a multi-round peer-battle by answering the seed question individually, criticizing the opponent’s weaknesses, and raising targeted follow-up queries to challenge the opponent further. During the multi-round battle process, the LLM’s true capabilities are drawn out and performance gaps become more visible. Lastly, a committee of LLM judge agents collectively discusses and evaluates the ability of the two candidates, mimicking the human voting process. By automating the entire evaluation process with LLM agents, we alleviate data contamination concerns in static datasets, reduce single-model bias with collective decision-making, and avoid long wait times for new models entering the human voting platform.
@@
To verify the evaluation framework, we run an extensive experiment with 17 LLMs. Compared to static and model-based benchmarks, Auto-Arena increases the Spearman correlation with human preferences by 4.5%, resulting in state-of-the-art alignment. Before and after peer battles, the Spearman correlation with human preferences increases by 46.4%, verifying our hypothesis that the peer battles can better display performance gaps. Before and after committee discussions, committee agreement increases by 20%, showcasing the effectiveness of the committee discussion mechanism. By studying the peer battles, we also discover intriguing LLM agent behaviors such as self-improvement and competitive actions. Using Chinese as an example, we further show that such an automated framework is easily extendable to non-mainstream languages and domains. Along with the project, we release a leaderboard, a website with demos, and the codebase. We are actively maintaining the leaderboard and altering the questions to provide reliable results in a timely manner.
2 Related Work
As LLMs evolve quickly, deriving trustworthy evaluations of their capabilities has become a challenge. Current evaluation methods can be divided into manual evaluation [5] and automatic evaluations. We primarily focus on automatic evaluations as they deliver more timely feedback. Automatic evaluations mainly consist of static datasets with predefined metrics and model-based metrics. Static datasets with predefined metrics [22, 6] suffer from contamination concerns and inflexibility in evaluating open-ended responses. Among model-based metrics, MT-Bench [44] and AlpacaEval [19] use LLM-as-a-judge to ask GPT-4 to compare model responses to a static dataset of questions. The model’s judgments achieve over 80% agreement with human preferences, proving the usability of using LLMs to evaluate response quality. Language-Model-as-an-Examiner [5] asks an LM examiner to construct knowledge-intensive questions within its memory, interact with the candidate in a series of follow-up queries, and rate the responses on dimensions including accuracy and factuality. KIEval [41] also incorporates an LLM-powered “interactor” role to examine deep comprehension of knowledge, which is shown to mitigate contamination issues on static datasets. However, such single-judge evaluations require the examiner to interact with each candidate parallelly, creating computational overheads and limiting the scope of queries. They also suffer from single-model bias, including bias towards LLM-generated summaries [28], inflated scores in multilingual evaluation [21], verbosity bias [19], and difficulties when evaluating candidates with close performance [36]. Therefore, some attempt to employ multiple judges agents. DRPE [39] uses multi-roleplayer prompting to mimic different roles with the same LLM and integrate outputs as votes for the final results. ChatEval [9] simulates different evaluator personas with the same base model to engage in debates, reaching a final evaluation result. While these methods come from the same motivation as our committee discussion component, they do not mitigate the model-specific bias as they use the same base LLM to simulate different personas. To further incorporate different LLMs as evaluators, PRD [25] allows 2 LLMs to discuss an evaluation and assigns higher voting weights to the LLM reviewers with stronger capabilities. Peer-review-in-LLMs [31] optimizes the voting weights as a learnable parameter. WideDeep [42] sets up a multi-layer neural network structure where each LLM evaluator functions as neurons. PRE [12] selects a small group of reviewers to produce evaluations individually and then asks a “chair” to produce the final evaluation using the aggregated evaluations. They show that the multi-agent approach effectively mitigates single-model bias. This line of work is similar to our “LLM judge committee” component. However, they are still limited to static datasets and specific domains.
Besides LLM evaluations, some works study competitive behaviors in LLMs, which is relevant to the LLM peer-battles in Auto-Arena. LM vs LM [14] shows that LLM cross-examinations can effectively discover factual errors. Debate [18] shows that multi-agent debate can improve factuality and reasoning. In MAD [27], LLM-debate can encourage divergent thinking, which helps tasks that require deep levels of contemplation. Khan et al. [24] shows that even non-expert weak LLMs can supervise expert LLMs if we allow the two LLM experts to engage in debates. Moreover, Zhao et al. [43] and Gu et al. [20] show interesting case studies where LLMs are engaged in simulated competitive environments and demonstrate human-like strategies.
Figure 2: The process of a Lincoln-Douglas-style peer-battle with the actions used. The <Think> action can be used by the candidates freely and is only visible to the candidate itself.
3 The Auto Arena Framework
@@
The Auto-Arena framework consists of three stages: question generation, peer battles, and committee discussions. These three stages are run sequentially and fully automated with LLM-based agents. Figure 1 shows the overall design. All prompts are included in Appendix A.1.
@@ 3.1 Question Generation
For debate questions, as using a static dataset could incur data contamination concerns and result in unfair evaluations, we ask an LLM examiner agent to dynamically generate questions. The examiner agent could be any capable LLM. Similar to MT-Bench [44], the generated questions cover 8 common categories in real-life conversations: writing, roleplay, extraction, reasoning, math, coding, STEM knowledge, and humanities/social science knowledge. In the prompt, the examiner is provided with a sample question and encouraged to generate diverse and difficult questions to ensure the depth and width of the debates examined. Examples of the generated questions is shown in Appendix A.2.
Specifically, as the examiner agent will also participate in the following debates, previous methods [5] could incur self-enhancement bias as they ask the examiner agents to only devise questions that they are confident about. We try to avoid such bias with two designs: 1. We do not disclose to the examiner that it will participate in this tournament. 2. We do not ask the examiner to only generate questions that it can solve. Thus, the questions do not cause much enhancement bias for the examiner.
@@ 3.2 Peer Debate
After the questions are drafted, we conduct peer battles around these questions among the candidate LLMs. In one peer battle, two candidate LLMs (A and B) debate around the given question, point out the opponent’s weaknesses, and raise follow-up questions that the opponent may fail to respond to.
In the peer battle, each candidate LLM can use 4 types of actions: 1. Think, where the candidate thinks about the question and plans its strategy. This action is hidden to the opponent. 2. Respond, where the candidate responds to the aforementioned question. 3. Criticize, where the candidate points out the loopholes and mistakes in the opponent’s previous responses. 4. Raise, where the candidate raises follow-up questions that are specifically designed to make the opponent expose its weaknesses.
The peer battle is designed according to the Lincoln-Douglas debate format  [^3] https://en.wikipedia.org/wiki/Lincoln-Douglas_debate_format, which is a competitive one-to-one debate used for popular competitions such as National Speech and Debate Association competitions. Overall, the peer battle consists of 3 rounds, where the candidates take turns to speak. The entire dialogue history is visible to both candidates. The process is illustrated in Figure 2. In the first round, A gives an initial response to the examiner’s question; B criticizes the weaknesses in A’s response and raises a targeted follow-up question; and A responds to B’s question. The second round has the same structure, with A and B roles reversed. In the third round, A and B cross-examine each other: A starts by criticizing B’s previous loopholes and raises follow-up questions. After responding, B also criticizes A’s loopholes and raises questions. A concludes the battle by responding again. In this process, both A and B get an equal number of each action to ensure fairness. To further reduce position bias, A and B’s order is randomly determined at the beginning of each debate.
Depending on which turn it is, we provide an action guide to the candidate, specifying the objectives and corresponding actions for this turn. Similar to human debate competitions, we time the candidates by providing a maximum length, which is also specified in the prompts. Any responses beyond the required length will be cut off. This design mitigates verbosity bias in LLM-as-a-judge [44], where LLM judges prefer longer and more verbose responses.
@@ 3.3 Committee Discussions
After the peer battle takes place, a committee of LLM judge agents will collectively determine the winner. The committee is always selected as the best five LLMs (except the participants) according to the current ranking. In the first round, the committee is initialized with MMLU [22] scores to approximate LLM performances. Judges are first asked to read through the battle history, elaborate judgment reasons, and give a decision on whether A is better, or B is better, or if there is a tie.
After the initial judgments are formed, the committee engages in a discussion. Each judge reads the other judge’s verdicts, decides whether to adjust the ratings, and elaborate the reasons. The collective intelligence introduces diverse viewpoints, improves judgment quality, and mitigates single-model bias. Finally, the winning candidate is decided by the majority voting of the discussed judgments.
For logical-reasoning questions that have ground-truth answers (reasoning, code, math), LLM-as-a-judge is known to show weak performances in judging the quality of responses [44]. We adopt prior approaches to establish the reference-based judge [44]. Specifically, we ask the strongest committee member to generate a reference answer and provide it to the judge when evaluating these responses.
@@ 4.1 Experimental Setup
For the main experiment, we first select 7 best or latest models from each popular model family on the top 30 list on Chatbot Arena with more than 10k votes each. To construct a comprehensive leaderboard [^4] https://huggingface.co/spaces/Auto-Arena/Leaderboard, we further add 10 models. Appendix A.4 provides a detailed list.
Baselines:
@@
For the baselines, we consider widely used evaluation benchmarks, including static datasets with fixed metrics and model-based metrics. A comparison table is shown in Appendix A.5.
@@
1. Static datasets with fixed metrics: (1) MMLU (Massive Multitask Language Understanding) [22], an extensive benchmark that covers 57 subjects and tests both world knowledge and problem-solving ability; (2) OpenLLM Leaderboard [6], a benchmark averaging performance metrics on 6 datasets.
@@
2. Static datasets with model-based metrics: (1) MT-Bench [44], a set of 80 multi-turn questions. Model responses are graded by GPT-4; (2) MT-Bench Hard [26], a benchmark dataset with 1,000 challenging user queries collected on Chatbot Arena. Model responses are graded by GPT-4; (3) Length-Controlled AlpacaEval [19], a benchmark based on AlpacaFarm evaluation set, which tests models’ abilities to follow general user instructions. Model responses are graded by GPT-4.
Setup:
@@
Among the 7 participants, we conduct a swiss-style tournament: For $n$ participants, instead of pairing each participant with $(n-1)$ others, a swiss-tournament pairs each player with $\lceil log_{2}(n)\rceil$ players of similar rankings without repeats. This design effectively reduces computational costs of ranking $n$ models from $O(n^{2})$ to $O(nlog_{2}(n))$. A cost analysis is included in Appendix A.5.
@@
Each candidate pair engages in 40 peer battles, with 5 GPT-4-generated questions from each of the 8 categories. As each battle consists of 3 rounds (each candidate speaks for 4 times), the competition scale is approximately the same as MT-Bench (80 questions, each candidate speaks twice). In the tournament, the rating scores are calculated with the ELO rating system [4, 7], which has become the standard practice in competitive games such as chess. Similar to the Chatbot Arena score calculation procedure [11], we compute the Bradley-Terry (BT) coefficients [8] for better statistical estimation.
@@
We initialize the Swiss tournament rankings according to MMLU scores, which is a static approximation of model performances. At the end of each pairing, we re-calculate ELO scores of current models. The committee is selected to be the remaining 5 models (besides the candidates). When the participant number increases, the committee is selected as the best 5 LLMs based on current ELO rankings. After the initial judgments, the committee members engage in one round of discussion.

### v4

# Auto-Arena:Automating LLM Evaluations with Agent Peer Battles and Committee Discussions

@@ Abstract
As LLMs continuously evolve, there is an urgent need for a reliable evaluation method that delivers trustworthy results promptly. Currently, static benchmarks suffer from inflexibility and unreliability, leading users to prefer human voting platforms like Chatbot Arena. However, human evaluations require significant manual effort. To address this, we propose the Auto-Arena, an innovative framework that automates the entire evaluation process using LLM-powered agents. Firstly, an LLM examiner generates questions. Then, two LLM candidates engage in a multi-round peer battle based on individual questions, aiming at revealing their true performance differences. Finally, a committee of LLM judges collaboratively discusses and decides the winner, reducing bias and enhancing fairness. During the peer battles, we observe intriguing scenarios where the LLM candidates display competitive behaviors and even learn from the opponents. In our extensive experiments involving 15 recent LLMs, Auto-Arena shows a 92.14% correlation with human preferences, surpassing all previous expert-annotated benchmarks without any manual efforts. As a result, Auto-Arena offers a promising alternative to current human evaluation platforms for evaluating LLMs automatically.
@@ 1 Introduction
Since ChatGPT and GPT-4 (OpenAI et al., 2024) gained popularity, Large Language Models (LLMs) have risen to the forefront of technological innovation, capturing broad industry and social interests (Wu et al., 2023b). This enthusiasm has spurred numerous organizations to release their own LLMs (Touvron et al., 2023; Team et al., 2024b). However, the rapid pace at which these models are released and updated poses a significant challenge for users attempting to understand their capabilities and monitor their evolution. Consequently, there has been a pressing demand for comprehensively evaluating LLMs recently (Chang et al., 2024a).
@@
The most popular existing method is automatic evaluation with static datasets. Among these, static datasets with predefined metrics, such as GSM8k (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021a), are constructed with aspect-specific input-output pairs, such as human exam-type questions and their corresponding answers. Given the questions, the LLM-produced answers are compared to ground-truth answers using metrics such as accuracy. This approach could suffer from inflexibility, contamination, and high human annotation costs. Firstly, the closed-form ground-truth answers limit their utility in assessing models’ performances on general or open-ended questions, which are the main use cases of LLMs. As the questions are static, they also risk contamination (Ravaut et al., 2024), where models may have been inadvertently exposed to elements of the test datasets during training, thereby skewing the evaluation results. The manual dataset construction also incurs high costs, creating barriers for extending to other domains or languages. As an alternative, static datasets with model-based evaluation, such as MT-Bench (Zheng et al., 2023) and AlpacaEval (Dubois et al., 2024a), evaluates LLMs on open-ended generations. These methods typically ask two models to generate responses to the same open-ended question and then employ a strong judge model (e.g., GPT-4) to choose the better response. However, the static question sets still bear contamination risks. Additionally, the assumption of the existence of a strong judge model makes the evaluation framework less generalizable and introduces model-specific bias.
Table 1: Comparison between Auto-Arena and other benchmarks or evaluation methods.
Questions Responses Judges
Method Dynamic? Auto-generated? Multi-turn? Open-ended? Auto? Committee?
OpenLLM Leaderboard
MMLU
GPQA
LC-AlpacaEval
MT-Bench
Arena-Hard
Chatbot Arena
Auto-Arena
@@
Aside from automated evaluations, human assessment, although requiring significant manual efforts, remains the gold standard for users. A notable example is Chatbot Arena (Zheng et al., 2023), a crowdsourcing platform that gathers anonymous votes on LLM performances and calculates Elo scores (Elo & Sloan, 1978) to rank these models. The resulting leaderboard [^1] https://leaderboard.lmsys.org/ is widely considered as a trustworthy indicator of LLMs’ general capabilities. However, a reliable model evaluation on this platform must be supported by a large number of human votes, which requires considerable time and effort. Consequently, when newly developed models enter the scene, they often struggle to quickly amass a large number of votes. Moreover, this strong reliance on human votes limits its application in various scenarios. For example, the performance of non-English languages is difficult to estimate, as most queries on the platform are in English. Moreover, the queries are mostly one-round and simple. The completely open participation may also result in uneven evaluation quality.
@@
To enable the evaluation of LLMs that is both automated and reliable while aligning with human preferences, we introduce Auto-Arena, a framework that automates the entire LLM evaluation process with LLM-powered agents. The framework consists of three stages: Firstly, an LLM examiner agent is tasked with generating questions, mimicking real-life users posting queries. Secondly, two LLM candidates interact with each other and engage in a multi-round peer battle by answering the seed question individually, criticizing the opponent’s weaknesses, and raising targeted follow-up queries to challenge the opponent further. During the multi-round battle process, the LLM’s true capabilities are drawn out and performance gaps become more visible. Lastly, a committee of LLM judges collectively discusses and evaluates the ability of the two candidates, mimicking the human voting process. As shown in Table 1, Auto-Arena has several key advantages compared to previous evaluation methods: First and foremost, instead of the simple and one-round question-answering scheme, Auto-Arena introduces a dynamic multi-round peer battle, which displays deeper abilities of LLMs, such as reasoning, interacting, and strategizing. The dynamic nature of peer battles also reduces contamination risks. Secondly, by expanding a single LLM judge into a committee of LLM judges, Auto-Arena alleviates potential model-specific evaluation bias. Finally, since the process of generating questions and judgments is fully automated in an end-to-end way, Auto-Arena can provide timely evaluations for new models and can easily extend to various domains and languages.
@@
To verify the reliability and alignment of the evaluation framework, we run an extensive experiment with 15 LLMs. Compared to static and model-based benchmarks, Auto-Arena results in the state-of-the-art alignment by achieving a 92.14% Spearman correlation with human preferences, surpassing all previous benchmarks. Although no manual efforts is involved, the high alignment with human preferences could originate from the human-like evaluation process, which is simulated using LLM agents. The extensive ablation experiments also demonstrate the reliability of the framework: Before and after peer battles, the Spearman correlation with human preferences increases by 5%, verifying our hypothesis that the peer battles can better display performance gaps. Before and after committee discussions, committee agreement increases by 11%, showing human-level agreement and verifying the effectiveness of the committee discussion mechanism. By studying the peer battles, we also discover intriguing LLM agent behaviors such as competitive and self-improvement actions. As the entire process is automatic, the evaluation can be easily adapted to other languages or domains by altering the prompts. We provide Chinese as a case study for extending to other languages.
In conclusion, our contributions can be summarized as follows:
  1. 1.
    We propose Auto-Arena, a fully automatic LLM evaluation framework where the examiner, candidates, and judges are all simulated with LLM-powered agents;
  2. 2.
    Specifically, we innovatively utilize peer battles for LLM evaluation, where two LLM agents engage in a multi-round debate. This process draws out the model’s deeper capabilities;
  3. 3.
    In our extensive experiment with 15 LLMs, we observe the state-of-the-art alignment with human preferences without any manual efforts;
  4. 4.
    During peer battles, LLM agents display intriguing behaviors, such as strategizing and learning from the opponents, which opens up possibilities for future work.
Figure 1: An illustration of Auto-Arena.
2 The Auto-Arena Framework
@@
As illustrated in Figure 1, the Auto-Arena framework consists of three stages: Question Generation, Multi-round Peer Battles, and Committee Discussions. These three stages are run sequentially and fully simulated with LLM-powered agents. All prompts are included in Appendix A.
@@ 2.1 Question Generation
For debate questions, as using a static dataset could incur data contamination concerns and result in unfair evaluations, we ask an LLM examiner agent to dynamically generate questions. The examiner agent could be any capable LLM. Similar to MT-Bench (Zheng et al., 2023), the generated questions cover 8 common categories in real-life conversations: writing, roleplay, extraction, reasoning, math, coding, STEM knowledge, and humanities/social science knowledge. The examiner is provided with a sample question and encouraged to generate diverse and difficult questions to ensure the depth and width of the evaluated debates. Examples of the generated questions are shown in Appendix B.
Specifically, as the examiner agent will also participate in the following debates, we try to alleviate self-enhancement bias with two designs: 1. We do not disclose to the examiner that it will participate in this tournament. 2. Previous methods (Bai et al., 2024) could incur self-enhancement bias as they ask the examiner agents to only devise questions that they are confident about. In comparison, we do not ask the examiner to only generate questions that it can solve. To further show that limited self-enhancement bias is present, we include an ablation study in Appendix E.
Figure 2: The process of a Lincoln-Douglas-style peer battle with the actions used. The <Think> action can be used by the candidates freely and is only visible to the candidate itself.
@@ 2.2 Peer Debate
After question generation, we conduct peer battles around these questions among the LLM candidates. In one peer battle, two LLM candidates (A and B) debate around the given question, point out the opponent’s weaknesses, and devise follow-up questions to further probe the opponent’s weaknesses.
In the peer battle, each candidate LLM has four available types of actions:
  • <Think>: The candidate generates internal thoughts about the question or plans a strategy. This action can be used at any time and remains concealed from the opponent.
  • <Respond>: The candidate answers the given question.
  • <Criticize>: The candidate identifies flaws and errors in opponent’s previous responses.
  • <Raise>: The candidate poses follow-up questions to reveal the opponent’s weaknesses.
The workflow of a peer battle takes the form of the Lincoln-Douglas debate format [^2] https://en.wikipedia.org/wiki/Lincoln-Douglas_debate_format. To help users better understand this debate format, we show the debate samples at https://auto-chatbot-arena.streamlit.app/., the most widely used one-on-one debate style in competitions such as those held by the National Speech and Debate Association. The peer battle consists of three rounds in which two candidate models alternate speaking. Both candidates can see the complete dialogue history. This process is depicted in Figure 2. In the first round, model A responds to the examiner’s initial question; model B criticizes the flaws in A’s response and raises a specific follow-up question; model A then responds to B’s follow-up question. The second round follows the same format, with A and B switching roles. In the third round, A and B cross-examine each other, starting with A criticizing the loopholes in B’s earlier responses and raising follow-up questions. After responding, model B criticizes A’s weaknesses and raises additional questions. Model A wraps up by responding once more. Throughout this process, both A and B perform an equal number of actions to maintain fairness. To minimize positional bias, the order of A and B is randomized at the start of each debate.
During the debate process, enhancement bias and contamination concerns are further reduced: The process of candidates raising follow-up questions to each other essentially decentralizes the question-generation process, reducing enhancement bias in the generated initial questions. Moreover, debating ensures that candidates are evaluated not only on their response to the initial question, but also in more comprehensive and deeper abilities, such as strategizing, criticizing the opponent, and drafting questions. In other words, answering the initial question well does not necessarily win the whole debate, which further reduce contamination concerns.
Depending on which turn it is, we provide an action guide to the candidate, specifying the objectives and corresponding actions for this turn. Similar to human debate competitions, we time the candidates by imposing a maximum length constraint, which is also specified in the prompts. Any responses beyond the required length will be cut off. This design mitigates verbosity bias in LLM-as-a-judge (Zheng et al., 2023), where LLM judges prefer longer and more verbose responses.
@@ 2.3 Committee Discussions
After the peer battle takes place, a committee of LLM judges collectively determines the winner. The committee is always selected as the five best LLMs according to the current ranking. To reduce bias, we exclude the participants themselves and models from the same family as the participants from the committee. For example, GPT-4 will not serve as a judge in evaluating a debate participated by GPT-3.5. In the first round, the committee is initialized with MMLU (Hendrycks et al., 2021a) scores to approximate LLM performances. Each judge is individually asked to read the entire peer battle history, elaborate judgment reasons, and give a decision on whether A is better, or B is better, or if there is a tie based on factors such as helpfulness, relevance, and accuracy.
After the initial judgments are formed, the committee engages in a discussion. In a discussion round, each judge reads the other judge’s verdicts in the previous rounds, elaborates its own thoughts for judgments, and drafts a discussed verdict. During the process, the judge may decide to adjust or maintain the previous judgments. Compared to the peer battles that exemplify multi-agent competitions, this committee discussion component synthesizes a multi-agent collaboration scheme. By enabling interactions among the judge agents and exchanges of different viewpoints, the discussion allows the committee to form a collective intelligence. As a result, it improves the judgment quality, boosts inter-judge agreement, and mitigates single-model bias. Finally, the winning candidate is decided by majority voting of the discussed judgments.
@@ 3.1 Experimental Setup
Model Selection:  For the main experiment, we first select 9 best or latest models that are representative of each popular model family on the top 30 list on the Chatbot Arena platform with more than 10k votes each at the time of experiments: GPT-4-0409-Turbo, GPT-3.5-Turbo-0125, Claude-3-Haiku, Qwen1.5-72B-Chat, Command-R+, Llama-2-70B-Chat, Mixtral-8x7b-Instruct-v0.1, Yi-34B-Chat, and Deepseek-LLM-67B. To construct a leaderboard, we further add 6 models that are newly released: GPT-4o-2024-05-13, Claude-3.5-Sonnet, Qwen2-72B-Instruct, Llama-3-70B, Gemma-2-27B, and Gemini-1.5-Flash. Appendix G provides a detailed list of the selected models.
@@
Baselines:   For the baselines, we consider popular evaluation benchmarks, including fixed metrics and model-based metrics. A comparison table is shown in Appendix H.
@@
1. Static datasets with fixed metrics: (1) OpenLLM Leaderboard (Beeching et al., 2023), a popular benchmark for open-source models averaging performance metrics on 6 key benchmarks, covering a large number of different evaluation tasks; (2) GPQA (Rein et al., 2023), a graduate-level google-proof Q&A benchmark consisting of 448 domain-expert-written questions written in scientific subjects; (3) MMLU (Massive Multitask Language Understanding) (Hendrycks et al., 2021a), an extensive benchmark that covers 57 subjects and tests both world knowledge and problem-solving ability;
@@
2. Static datasets with model-based metrics: (1) MT-Bench (Zheng et al., 2023), a set of 80 multi-turn questions. Model responses are graded by GPT-4; (2) Arena Hard (Li* et al., 2024), a benchmark dataset with 1,000 challenging user queries collected on Chatbot Arena. Model responses are graded by GPT-4-Turbo; (3) Length-Controlled AlpacaEval (Dubois et al., 2024a), a benchmark based on AlpacaFarm evaluation set (Dubois et al., 2024b), which tests models’ abilities to follow general user instructions. Models are evaluated by their win rates against GPT-4-Turbo, graded by GPT-4-Turbo.
@@
Setup:   Among the 9 participants, we conduct a swiss-style tournament: For $n$ participants, instead of pairing each participant with $(n-1)$ others, a swiss-tournament pairs each player with $\lceil log_{2}(n)\rceil$ players of similar rankings without repeats. This design effectively reduces computational costs of ranking $n$ models from $O(n^{2})$ to $O(nlog_{2}(n))$. A cost analysis is included in Appendix H.
@@
Each candidate pair engages in 40 peer battles, with 5 questions from each of the 8 task categories that are specified in Section 2.1. We provide studies showing that the generated questions can reduce contamination concerns in Appendix C and are generalizable to real-world scenarios in Appendix D. As each battle consists of 3 rounds (each candidate speaks for 4 times), the competition scale is approximately the same as MT-Bench (80 questions, each candidate speaks twice). In the tournament, the rating scores are calculated with the Elo rating system (Bai et al., 2022; Boubdir et al., 2023), which has become the standard practice in competitive games such as chess (Elo & Sloan, 1978). Similar to the Chatbot Arena score calculation procedure (Chiang et al., 2024), we compute the Bradley-Terry (BT) coefficients (Bradley & Terry, 1952) for better statistical estimation. Following the Reference-Guided judge in Zheng et al. (2023), we ask the best-performing judge to give a reference answer for evaluating logical-reasoning questions (math, coding, reasoning).
@@
We initialize the Swiss tournament rankings according to MMLU scores, which is a static approximation of model performances. At the end of each pairing, we re-calculate Elo scores of current models. The committee is selected as the best 5 LLMs based on current Elo rankings at each round. After forming initial judgments, the committee members engage in one round of discussion. The final result is decided by majority voting of the discussed judgments.
Table 2: Correlations with Chatbot Arena Elos of evaluation benchmarks on 9 LLMs.
Spearman
Correlation
OpenLLM (Beeching et al., 2023) -15.39%
GPQA (Rein et al., 2023) 36.84%
MMLU (Hendrycks et al., 2021b) 56.36%
LC-AlpacaEval (Dubois et al., 2024a) 82.14%
MT-Bench (Zheng et al., 2023) 82.86%
Arena-Hard (Li* et al., 2024) 85.71%
Auto-Arena 91.67%
w/o Peer Battles 86.67%
w/o Committee Discussions 88.33%